What is the 17th term in the arithmetic sequence in which a^6 is 101 and a^9 is 83?
Answer by richwmiller(17219)

an = a + (n - 1)
101 = a + 5d,
83= a + 8d,
a17 = a + 16d
a17=131-6*(16)
a17=131-96
a17=35

a = 131, d = -6, t = 35